# Arrange, Inpaint, and Refine: Steerable Long-term Music Audio Generation and Editing via Content-based Controls

### v1

@@ 4.1 Dataset
Our model is trained on the training split of Slakh2100, a synthetic music audio dataset with MIDI annotations Manilow et al. (2019). The total training set comprises 1289 instrumental songs with silenced onsets and tails removed. Evaluation is conducted on both the test split (151 songs) of Slakh2100 and RWC-POP100, a dataset featuring 100 real-recording pop songs with MIDI, chord, and beat annotations Goto et al. (2002).
@@ 4.2 Training
We train three models: Drums-AIR, Chord-AIR, and Piano-AIR, tailored for drum-track conditioning, chord-progression conditioning, and piano-cover conditioning, respectively.
4.2.1 Configuration
@@
[…]
Full Prefix Full Prefix
  Drum-AIR 0.749 0.756 1.423 1.422
Chord-AIR 0.753 0.757 1.220 1.222
Piano-AIR 0.755 0.761 1.290 1.282
  MusicGen 0.656 0.687 1.251 1.218
VampNet 0.631 0.643 2.910 3.424
(a) Slakh2100 Test Set
[…]
Full Prefix Full Prefix
  Drum-AIR 0.619 0.627 1.606 1.691
Chord-AIR 0.614 0.625 1.593 1.681
Piano-AIR 0.611 0.621 1.531 1.623
  MusicGen 0.373 0.441 2.474 2.276
VampNet 0.613 0.618 3.689 3.910
(b) RWC-POP-100
Table 1: Evaluation of the models in terms of inpainting musicality. Adapter width $m=50$.
[…]
  Drum-AIR 6.676 - -
Chord-AIR - 0.745 0.830
Piano-AIR - 0.720 0.849
(a) Slakh2100 Test Set
[…]
  Drum-AIR 6.531 - -
Chord-AIR - 0.650 0.837
Piano-AIR - 0.615 0.845
(b) RWC-POP-100
Table 2: Evaluation of the models in terms of controllability. Adapter width $m=50$.
(a) Drum track controls, where the condition is drum audio.
(b) Chord progression controls, where the condition is block chords audio.
(c) Arrangement and orchestration from piano cover, where the condition is piano cover audio.
Figure 4: Inpainting with various controls. The spectrogram conceptually repeats the same music segment twice, with the left half ($1,...,T$) providing the model with the infilling context and the controls, and the right half ($T+1,...,2T$) for the model to predict. Specifically, in the left half, outside the blue rectangle is the unmasked infilling context, and inside the blue rectangle is the separated/synthesized/user-specified audio representation of the condition (e.g., (a) drum, (b) block chords, (c) piano cover). In the right half, the orange rectangle highlights the inpainted results.
4.2.2 Masking Scheme
Throughout the training process, we uniformly sample a mask ratio $r$ from the interval $[0.4,0.8]$ and generate a binary mask with this ratio. To maintain the continuity of audio segments, we utilize a median filter with a window size of 11 to smooth the mask before applying it to the input Encodec tokens.
@@ 4.3 Evaluation
We assess the performance of our models using two distinct datasets: the synthetic instrumental dataset (Slakh2100 Test Set) and a real-recording vocal dataset (RWC-POP-100). From each test song, we extract a 15-second audio segment for evaluation. We present three types of masks to corrupt the test audio, resulting in discontinuous masking areas: one, two, and four masking areas, respectively, referred to as Mask 1, Mask 2, and Mask 3, as shown in Figure 5. The total duration of masking areas sums up to 7.5s for each setting. The metrics we employ are as follows:
@@
    Source-to-Distortion Ratio (SDR). For Drum-AIR, we quantify the SDR between the inpainted drum tracks (separated by Demucs) and the condition drum tracks, following newest SDR definition from the Music Demixing Challenge (MDX 2021) competition Mitsufuji et al. (2022).
@@
    Chord Accuracy. For both Chord-AIR and Piano-AIR, we compare the inpainted chords (predicted by a chord recognition model) with the condition chord progression. We report weighted recall scores with a time resolution of 20ms for chord accuracy.
@@
    Chroma Cosine Similarity. For both Chord-AIR and Piano-AIR, we evaluate the chroma cosine similarity between the inpainted audios and the condition tracks, with a window size of 2048 and a hop size of 640 for the chromagram.
@@
    CLAP Score. For each model, we report the CLAP score Wu et al. (2023), assessing the semantic distance between the inpainted audio and the original audio.
@@
    Fréchet Audio Distance (FAD). Within each test set, we calculate the FAD of VGG (Visual Geometry Group) features between the inpainted audio and the original audio.
@@ 4.4 Baselines
We designate MusicGen Copet et al. (2023) and VampNet Garcia et al. (2023) as our baselines, as there are currently no other comparable content-based controllable inpainting models available for comparison. As MusicGen can only perform continuation rather than inpainting, we only evaluate MusicGen using the past context of Mask 1.
@@ 4.5 Results
As illustrated in Table 2 and 1, our model demonstrates competitive performance compared to unconditional generation and inpainting models. Furthermore, it showcases promising steerabilty. More audio demos are available here. [^6] https://kikyo-16.github.io/AIR/##more.
4.5.1 Long-Gap Inpainting
Table 1 presents the semantic similarity between inpainted audios and the original recordings, along with the audio qualities. In this scenario, MusicGen receives the prefix of Mask 1 (3.75s) and is tasked with generating an 11.25-second continuation. Conversely, VampNet and AIR models are conditioned with both the prefix and suffix of Mask 1. We present metrics for prefix-CLAP and prefix-FAD, evaluating the predicted (repeated) prefix and inpainted segments (11.25s), while full-CLAP and full-FAD assess both the predicted prefix, the inpainted segment, and the predicted suffix.
The comparison between prefix-CLAP and full-CLAP values in Table 1(a) indicates that as MusicGen generates longer continuations, it “strays further” from the given context. When employing our heterogeneous adapter to enable inpainting settings for MusicGen, the resulting audio exhibits promising fidelity in preserving the original semantic context during generation. Furthermore, the low FAD further indicates that audio samples inpainted by our model exhibit competitive naturalness compared to those generated by the vallina MusicGen without suffix constraints. Compared to VampNet, our model achieves superior long-gap inpainting, as evidenced by the FAD value.
Table 1(b) presents the results on the real-recording dataset with vocals, which is Out-Of-Distribution (OOD) w.r.t. the training dataset of MusicGen and our models but not w.r.t. VampNet. Comparing these results with those in Table 1(a), it is evident that our models exhibit superior generalization ability, outperforming the vanilla MusicGen.
@@ 4.5.1 Long-Gap Inpainting
Table 3: The performance of models with different $m$ on Mask 2.
(a) $b^{l}_{1}$
(b) $b^{l}_{2}$
(c) $b^{l}_{3}$
(d) $b^{l}_{4}$
Figure 7: The variation of $|g^{l}_{t}|$ across different types of adapters during training.
4.5.2 Controllable Inpainting
As demonstrated in Table 1, our model excels in controllable inpainting across various music editing tasks. The high SDR values in Table 2 indicate our model’s proficiency in internal control, effectively inpainting the masked areas while preserving the condition track. Additionally, chord, chroma, and beat metrics underscore the capability of our models to integrate controls. Furthermore, CLAP values in Table 1 highlight that applying content-based controls extracted from the original audio serve as an auxiliary mechanism for maintaining the semantics of the original audio.
4.5.3 Arrangement and Refinement
Figure 4 illustrates examples of the results for the score-conditioned arrangement and the track-conditioned refinement tasks, highlighting our model’s capability to reconstruct missing segments of the spectrogram. This demonstrates our approach’s adaptability in performing spectrogram inpainting under various conditions.
More specifically, Figure 6 presents a case of inpainting where 70% of the spectrogram was intentionally masked prior to the inpainting process. Despite the significant loss of information, our model is able to restores the music composition, showing its robustness in handling substantial data gaps.
4.5.4 Low-Resource Fine-Tuning
Referring to Table 3, our heterogeneous adapters demonstrate a notable advantage in low-resource settings, with trainable parameters totaling less than 20M and a storage space requirement of less than 75M for the largest model. In comparison with the 3.3B parameters of vanilla MusicGen, our method effectively facilitates fine-tuning, converting an autoregressive music audio LLM into a steerable inpainting model.
@@ 4.6 Ablation
4.6.1 Impact of Different Adapter Size
A smaller $m$ implies a reduced size of adapters, resulting in fewer trainable parameters. However, Table 3 illustrates that a small $m$ does not significantly impair performance, highlighting the effectiveness of the proposed heterogeneous adapters. Additionally, Figure 7 displays the variation of absolute values of learnable gates $g^{l}_{t}$ in different layers during training. It reveals that the adapter influences deep layers more than shallow layers and that the adapters ${\bm{a}}^{l}_{\leq 2}$ for the prefix area encode more information than the prediction area ${\bm{a}}^{l}_{>2}$.
4.6.2 Impact of Different Masks
During the ablation study, we utilize three different masks, as illustrated in Figure 5, and report the FAD values of inpainted audios in Table 4. The context becomes more continuous, and the gap to be filled becomes shorter as we progress from Mask 1 to Mask 3. The results indicate that our model is highly robust to different mask structures and demonstrates a promising long-gap inpainting capability. Conversely, the baseline model’s performance deteriorates as the masked gap lengthens.
  Mask 1 Mask 2 Mask 3
  Drum-AIR 1.423 1.442 1.467
VampNet 2.910 2.390 2.091
(a) Slakh2100 Test Set
  Mask 1 Mask 2 Mask 3
  Drum-AIR 1.606 1.673 1.759
VampNet 3.689 3.496 3.336
(b) RWC-POP-100
Table 4: The FAD values of generated audio with different masked contexts. A lower value indicates higher audio quality.

### v2

@@ 4.1 Dataset
Our model is trained on the training split of Slakh2100, a dataset of synthesized MIDI files using a high-quality soundfont Manilow et al. (2019). We derive pseudo chord and beat annotations using Madmom Böck et al. (2016) and a chord recognition model Jiang et al. (2019). The total training set comprises 1289 instrumental songs with silence trimmed away at both ends. Evaluation is conducted on both the test split (151 songs) of Slakh2100 and RWC-POP100, a dataset featuring 100 real-recording pop songs with MIDI, chord, and beat annotations Goto et al. (2002).
@@ 4.2 Training
We train three models: Drum-AIR, Chord-AIR, and Piano-AIR, tailored for drum-track conditioning, chord-progression conditioning, and piano-cover conditioning, respectively.
@@ 4.3 Evaluation
Throughout the training process, we uniformly sample a mask ratio from the interval $[0.4,0.8]$ and generate a binary mask with this ratio. To maintain the continuity of audio segments, we utilize a median filter with a window size of 11 to smooth the mask before applying it to the input Encodec tokens.
We assess the performance of our models using two distinct datasets: the synthetic instrumental dataset (Slakh2100 Test Set) and a real-recording vocal dataset (RWC-POP-100). From each test song, we extract a 15-second audio segment for evaluation. We present three types of masks to corrupt the test audio, resulting in discontinuous masking areas: one, two, and four masking areas, respectively, referred to as Mask Pattern 1 through 3, as shown in Figure 7. The total duration of masking areas sums up to 7.5s for each setting.
The metrics we employ are as follows:
@@
    Source-to-Distortion Ratio (SDR). To evaluate how closely Drum-AIR adheres to the drum-track condition, we measure the SDR between the inpainted drum tracks (separated by Demucs) and the condition drum tracks. This measurement uses the latest SDR definition as specified in the Music Demixing Challenge (MDX 2021) competition Mitsufuji et al. (2022).
@@
    Chord Accuracy. For both Chord-AIR and Piano-AIR, we evaluate adherence to the condition chord progression by comparing the inpainted chords (predicted by a chord recognition model Jiang et al. (2019)) with the condition chords. We report the weighted recall scores with a time resolution of 20ms to measure chord accuracy.
@@
    Chroma Cosine Similarity. To assess how well Chord-AIR and Piano-AIR adhere to the chord and piano conditions, we calculate the chroma cosine similarity between the inpainted audios and the condition tracks, with a window size of 2048 and a hop size of 640 for the chromagram.
@@
    CLAP Score. For each model, we report the CLAP score Wu et al. (2023b), which assesses the semantic distance between the inpainted audio and the original audio.
@@
    Fréchet Audio Distance (FAD). For each model, we calculate the Frechet Audio Distance (FAD) of Visual Geometry Group (VGG) features between the inpainted audio and the original audio.
@@ 4.4 Baselines
We designate MusicGen Copet et al. (2023) and VampNet Garcia et al. (2023) as our baselines, as there are currently no other comparable content-based controllable inpainting models available for comparison. As MusicGen can only perform continuation rather than inpainting, we only evaluate MusicGen using the past context of Mask Pattern 1.
@@ 4.5 Results
As illustrated in Table LABEL:table:exp2 and Table LABEL:table:exp1, our model demonstrates competitive performance compared to unconditional generation and inpainting models. Furthermore, it showcases promising steerabilty.
Long-Gap Inpainting.
Table LABEL:table:exp2 presents the semantic similarity between inpainted audios and the original recordings, along with the audio qualities. In this scenario, MusicGen receives the prefix of Mask Pattern 1 (3.75s) and is tasked with generating an 11.25-second continuation. Conversely, VampNet and AIR models are conditioned with both the prefix and suffix of Mask Pattern 1. We present metrics for prefix-CLAP and prefix-FAD, evaluating the predicted (repeated) prefix and inpainted segments (11.25s), while full-CLAP and full-FAD assess both the predicted prefix, the inpainted segment, and the predicted suffix.
The comparison between prefix-CLAP and full-CLAP values in Table LABEL:table:exp2(a) indicates that as MusicGen generates longer continuations, it “strays further” from the given context. When employing our heterogeneous adapter to enable inpainting settings for MusicGen, the resulting audio exhibits promising fidelity in preserving the original semantic context during generation. Furthermore, the low FAD further indicates that audio samples inpainted by our model exhibit competitive naturalness compared to those generated by the vallina MusicGen without suffix constraints. Compared to VampNet, our model achieves superior long-gap inpainting, as evidenced by the FAD value.
Table LABEL:table:exp2(b) presents the results on the real-recording dataset with vocals, which is Out-Of-Distribution (OOD) w.r.t. the training dataset of MusicGen and our models but not w.r.t. VampNet. Comparing these results with those in Table LABEL:table:exp2(a), it is evident that our models exhibit superior generalization ability, outperforming the vanilla MusicGen.
@@
Table 2: The performance of models with different $m$ on Mask Pattern 2.
Controllable Inpainting.
As demonstrated in Table LABEL:fig:editing, our model excels in controllable inpainting across various music editing tasks. The high SDR values in Table LABEL:table:exp1 indicate our model’s proficiency in internal control, effectively inpainting the masked areas while preserving the condition track. Additionally, chord, chroma, and beat metrics underscore the capability of our models to integrate controls. Furthermore, CLAP values in Table LABEL:table:exp2 highlight that applying content-based controls extracted from the original audio serve as an auxiliary mechanism for maintaining the semantics of the original audio.
Arrangement and Refinement.
Figure LABEL:fig:spec1 illustrates examples of the results for the score-conditioned arrangement and the track-conditioned refinement tasks, highlighting our model’s capability to reconstruct missing segments of the spectrogram. This demonstrates our approach’s adaptability in performing spectrogram inpainting under various conditions.
Figure 10 presents a case of inpainting where 70% of the spectrogram was intentionally masked prior to the inpainting process. Despite the significant loss of information, our model is able to restore the music composition, showing its robustness in handling substantial data gaps.
Low-Resource Fine-Tuning.
As shown in Table 2, our heterogeneous adapters demonstrate a notable advantage in low-resource settings, with trainable parameters totaling less than 20M and a storage space requirement of less than 75M for the largest model. In comparison with the 3.3B parameters of vanilla MusicGen, our method effectively facilitates fine-tuning, converting an autoregressive music audio LLM into a steerable inpainting model.
@@ 4.6 Ablation
Impact of Different Adapter Size.
A smaller $m$ implies a reduced size of adapters, resulting in fewer trainable parameters. However, Table 2 illustrates that a small $m$ does not significantly impair performance, highlighting the effectiveness of the proposed heterogeneous adapters. Additionally, Figure LABEL:fig:exp:gates displays the variation of absolute values of learnable gates $g^{l}_{r(t)}$ in different layers during training. It reveals that the adapter influences deep layers more than shallow layers and that the adapters ${\bm{a}}^{l}_{\leq 2}$ for the prefix area encode more information than the prediction area ${\bm{a}}^{l}_{>2}$.
Impact of Different Mask Patterns.
During the ablation study, we utilize three different mask patterns, as illustrated in Figure 7, and report the FAD values of inpainted audios in Table LABEL:table:exp4. The context becomes more continuous, and the gap to be filled becomes shorter as we progress from Mask Pattern 1 to Mask Pattern 3. The results indicate that our model is highly robust to different mask structures and demonstrates a promising long-gap inpainting capability. Conversely, the baseline model’s performance deteriorates as the masked gap lengthens.
@@ 5 Conclusions
Acknowledgments
This work was supported in part through the NYU and NYUSH IT High Performance Computing resources, services, and staff expertise.